# appReferences in the Appendix

### v2

@@ Abstract
Online linear programming (OLP) has found broad applications in revenue management and resource allocation. State-of-the-art OLP algorithms achieve low regret by repeatedly solving linear programming (LP) subproblems that incorporate updated resource information. However, LP-based methods are computationally expensive and often inefficient for large-scale applications. In contrast, recent first-order OLP algorithms are more computationally efficient but typically suffer from worse regret guarantees. To address these shortcomings, we propose a new algorithm that combines the strengths of LP-based and first-order OLP methods. The algorithm re-solves the LP subproblems periodically at a predefined frequency $f$ and uses the latest dual prices to guide online decision-making. In addition, a first-order method runs in parallel during each interval between LP re-solves, smoothing resource consumption. Our algorithm achieves $\mathcal{O}(\log(T/f)+\sqrt{f})$ regret, delivering a “wait-less” online decision-making process that balances the computational efficiency of first-order methods and the superior regret guarantee of LP-based methods.
@@ 1 Introduction
Sequential decision-making has garnered increasing attention for its utility in guiding optimal strategies in dynamic environments. It addresses the problems of identifying the optimal decisions and policies in environments where knowledge of the system continuously accumulates and evolves over time. In this paper, we study Online Linear Programming (OLP) [1], a powerful paradigm that encapsulates the core principles of sequential decision-making. OLP methods have been extensively applied to diverse areas, including resource allocation [2], online advertising [19], and inventory management [21].
@@
In the OLP problem, customers arrive sequentially, each presenting a resource request and an associated bidding price. The objective is to determine which resource requests to fulfill to maximize total revenue while ensuring that no resource is ever over-allocated. The key challenge is that decisions must be made immediately and irrevocably, relying solely on historical data without knowledge of future arrivals. The goal is to minimize the cumulative regret with respect to the optimal hindsight linear programming (LP) solution.
@@
Most OLP algorithms use online learning methods to gradually refine their estimates of the optimal dual prices, using these values to guide near-optimal decision-making. Commonly, these algorithms rely either on LP-based or first-order methods. LP-based methods repeatedly solve linear programs at each timestep with all available information, thereby updating the dual prices. However, the substantial computational demands of LP-based methods often restrict their application in time-sensitive settings. In contrast, first-order methods leverage gradient information for quick, incremental updates to the dual prices. However, they generally cannot achieve the same $\mathcal{O}(\log{T})$ regret guarantee as the LP-based methods [16]. These trade-offs motivate an open question at the intersection of online learning and decision making:
Can we simultaneously achieve low regret and efficiency?
We answer this question affirmatively in the well-studied stochastic input model–where the resource request and bidding price for each customer are sampled identically and independently from an unknown distribution—by proposing a decision-making framework that combines LP-based and first-order OLP algorithms. Our algorithm iteratively optimizes the dual prices, which progressively converge to the optimal dual solution.
1.1 Main Contributions
We propose a parallel multi-stage framework for OLP that integrates online learning and decision-making. Building on this framework, we develop a mathematical analysis that reconciles existing OLP approaches, ultimately allowing us to present an algorithm that achieves logarithmic worst-case regret. We formalize the trade-off between decision quality and computational efficiency, enabling end-users to quantify how allocating additional computational resources will reduce regret. In more detail, our contributions can be summarized as follows:
@@ 1.1 Main Contributions
  • Parallel Multi-stage Framework: We separate online learning and decision-making into two separate yet interlinked processes, connected through a feedback loop. We re-solve the LP subproblems periodically at a fixed frequency and pass the result to the decision-making path. These results guide decision-making until the next time that the LP is re-solved. Additionally, to enhance the algorithm’s efficiency, we apply a first-order method during the initial and final stages, restarting with the most recent learning outcomes. By integrating the LP-based and first-order techniques, this approach leverages their respective strengths.
@@
  • Regret Analysis: We develop a mathematical analysis for our algorithm where we derive a new performance metric to account for the inter-dependency between the LP-based and first-order methods. We unify the analysis of these two methods and provide a “spectrum theorem” that bounds regret for any feasible choice of the re-solving frequency in $[1,T]$:
@@
    If we re-solve the updated linear programs every $f$ timesteps over the total time horizon of length $T$ and utilize gradient information in the initial and final $f$ steps, our algorithm is guaranteed to achieve a worst-case regret of
    $\mathcal{O}\big{(}\log\big{(}\tfrac{T}{f}\big{)}+\sqrt{f}\big{)}.$ (1)
    When $f=1$, the algorithm is a pure LP-based method and reduces to [16], and when $f=T$, the algorithm is a pure first-order method with regret $\mathcal{O}(\sqrt{T})$. In our analysis, one can choose the “optimal” $f$ to balance the LP-solving times and regret. Compared to [23], our method enables a “wait-less” decision-making system for customers during the whole process.
@@ 1.2 Related Literature
Online resource allocation and OLP have been widely studied under two predominant models: the stochastic input model [9, 6] and the stochastic permutation model [1, 11]. In our paper, we study the stochastic input model, where each customer’s resource request and bidding price—represented as columns of the constraint matrix—are drawn i.i.d. from an unknown distribution $\mathcal{P}\in\Xi$.
Table 1: Performances in current OLP literature
@@
[16] Bounded, continuous support, non-degeneracy LP-based $\mathcal{O}(\log T\log\log T)$ Delay
[4] Bounded, continuous support, non-degeneracy LP-based $\mathcal{O}(\log T)$ Delay
[17] Bounded, continuous support, non-degeneracy LP-based $\mathcal{O}(\log T)$ Delay
[12] Bounded, finite support LP-based $\mathcal{O}(\log^{2}T)$ Delay
[5] Bounded, finite support, non-degeneracy LP-based $\mathcal{O}(1)$ Delay
[14] Bounded, finite support, non-degeneracy LP-based $\mathcal{O}(1)$ Delay
[23] Bounded, continuous support, non-degeneracy LP-based $\mathcal{O}(\log T/f)$ Delay
@@
[15] Bounded First-order $\mathcal{O}(\sqrt{T})$ No Delay
[2] Bounded First-order $\mathcal{O}(\sqrt{T})$ No Delay
[7] Bounded First-order $\mathcal{O}(\sqrt{T})$ No Delay
[3] Bounded First-order $\mathcal{O}(\sqrt{T})$ No Delay
[20] Bounded, finite support, non-degeneracy First-order $\mathcal{O}(T^{3/8})$ No Delay
[8] Bounded, continuous support, non-degeneracy First-order $\mathcal{O}(T^{1/3})$ No Delay
[17] Bounded, continuous support, non-degeneracy First-order $\mathcal{O}(\log^{2}T)$ No Delay
LP-based Methods
This class of algorithms focuses on the OLP dual problem, repeatedly solving it to update the decision variables. While these methods achieve low regret, they are computationally expensive. The dual LPs are formulated based on the information revealed by each arriving customer. Early versions of these algorithms maintained a fixed average resource consumption rate throughout the time horizon [1], whereas more recent variants adaptively adjust the consumption rate based on the remaining resources, updating the LP accordingly [16]. Under standard non-degeneracy and continuous support assumptions, these LP-based methods can achieve a regret bound of $\mathcal{O}(\log T)$. Several related works explore variants of this approach: [4] studies a multi-secretary problem and also attains $\mathcal{O}(\log T)$ regret; [17] incorporates a regularization term on the resource and also achieves $\mathcal{O}(\log T)$ regret; [12] assumes that the resource distribution has a finite support and obtains $\mathcal{O}(\log^{2}T)$ regret; and [5] makes a stronger assumption that both the resource and price distributions have finite support, leading to a constant regret.
First-order Methods
These algorithms leverage gradient information to update the decision variables directly without explicitly solving LPs. Although less accurate than LP-based methods, they enable more efficient computation and are therefore well-suited for large-scale or time-sensitive applications. First-order methods can attain a regret of $\mathcal{O}(\sqrt{T})$ with mirror descent and subgradient information under distributions with continuous support [15, 2], and $\mathcal{O}(T^{3/8})$ under non-degeneracy assumptions for finite support distributions [20]. Furthermore, [7] attains a $\mathcal{O}(\sqrt{T})$ regret under proximal point updates, while [3] obtains the same regret via a momentum variant of mirror descent. [8] utilizes a restart strategy to decouple the learning and decision-making processes and improves the regret to $\mathcal{O}(T^{1/3})$. [17] integrates the resource consumption rate adjustment from LP-based methods into first-order methods and obtains $\mathcal{O}(\log^{2}T)$ regret.
Delay in Decision-making
This line of research examines the impact of decision delays in online resource allocation. [10] analyzes an online setting with a mix of impatient customers, whose requests must be decided upon immediately and irrevocably, and partially patient customers, for whom decisions can be postponed for several time periods. [22] demonstrates that deferring decisions—either by introducing explicit delays or batching customers—can substantially reduce regret. In concurrent research, [23] partially solves this delay issue by solving LPs in batches. While their approach is similar in spirit to ours, they make a stronger assumption that each resource request has a positive lower bound in expectation and their algorithm still requires customers to wait for decisions in the initial and final batches. Compared with the aforementioned works, our framework imposes minimal assumptions on the problem parameters, attains a regret bound of $\mathcal{O}(\log(T/f)+\sqrt{f})$, and provides a “wait-less” decision-making process throughout the entire time horizon.
The rest of the paper is organized as follows. Section 2 introduces the problem formulation, preliminary dual convergence results, and the main assumptions. Section 3 proposes our novel framework to integrate the LP-based and first-order methods and provides our “spectrum theorem,” which bounds the regret of our algorithm by $\mathcal{O}(\log(T/f)+\sqrt{f})$ for any choice of $f$. Section 4 presents experiments that illustrate the strong performance of our algorithm and validate our theoretical conclusions.
@@ 2 Problem Setup
[…]
2.1 Online LP Formulation
Consider an online resource allocation problem over the time horizon $T$. Initially, we have an inventory vector $\bm{b}\in\mathbb{R}^{m}$, representing $m$ types of resources. At each time step $t$, a customer arrives with a request characterized by $(r_{t},\bm{a}_{t})\sim\mathcal{P}$, where $r_{t}\in\mathbb{R}$ is the offered payment (bid), $\bm{a}_{t}\in\mathbb{R}^{m}$ is the customer’s demand for each resource, and $\mathcal{P}$ is a fixed, unknown distribution. We must decide whether to accept or reject this request, represented by the decision variable $x_{t}\in\{0,1\}$. The goal is to maximize the cumulative reward. This problem can be formulated as the following OLP, which we refer to as the primal linear program (PLP):
$\max_{\bm{x}}$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{t=1}^{T}r_{t}x_{t}$ (PLP)
subject to $\leavevmode\nobreak\ \quad\sum_{t=1}^{T}a_{it}x_{t}\leq b_{i},i=1,...,m$
[…]
@@ 2.1 Online LP Formulation
$\min_{\bm{p},\bm{y}}$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{i=1}^{m}b_{i}p_{i}+\sum_{t=1}^{T}y_{t}$ (DLP)
subject to $\leavevmode\nobreak\ \quad\sum_{i=1}^{m}a_{it}p_{i}+y_{t}\geq r_{t},t=1,...,T$
$\quad\ \ p_{i},y_{t}\geq 0,\forall i,t,$
where $\bm{p}$ is the vector of dual prices. Let $\bm{d}=\bm{b}/T\in\mathbb{R}^{m}$ denote the initial average resource capacity. As demonstrated by [15], (DLP) can be written as
$\min_{\bm{p}}$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ f_{T}(\bm{p})\coloneqq\frac{1}{T}\sum_{t=1}^{T}\left[\sum_{i=1}^{m}d_{i}p_{i}+\left(r_{t}-\sum_{i=1}^{m}a_{it}p_{i}\right)^{+}\right]$ (2)
subject to $\leavevmode\nobreak\ \quad p_{i}\geq 0,i=1,...,m$
This can be regarded as a $T$-sample approximation of the following stochastic program:
$\min_{\bm{p}}$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ f(\bm{p})\coloneqq\mathbb{E}[f_{T}(\bm{p})]=\bm{d}^{\top}\bm{p}+\mathbb{E}[(r-\bm{a^{\top}p})^{+}]$ (3)
subject to $\leavevmode\nobreak\ \quad\bm{p}\geq\bm{0}$
where the expectation is taken with respect to the coefficient pair $(r,\bm{a})$ drawn from the unknown distribution $\mathcal{P}$.
We define the optimal solutions to the $T$-sample approximation problem (2) and the stochastic program (3) respectively as
$\bm{p}_{T}^{*}=\arg\min_{\bm{p}\geq\textbf{0}}\leavevmode\nobreak\ f_{T}(\bm{p})\quad\text{and}\quad\bm{p}^{*}=\arg\min_{\bm{p}\geq\textbf{0}}\leavevmode\nobreak\ f(\bm{p}).$
The decision variables $x_{t}^{*}$ can be established from the complementary slackness condition as
$x_{t}^{*}=\left\{\begin{array}[]{ll}\ 0,&r_{t}<\bm{a}_{t}^{\top}\bm{p}_{T}^{*},\vspace{3 pt}\\ \ 1,&r_{t}>\bm{a}_{t}^{\top}\bm{p}_{T}^{*}.\end{array}\right.$ (6)
@@
This connection between the primal-dual optimal solutions motivates us to propose dual-based online LP algorithms and then provide a convergence analysis of $\bm{p}_{T}^{*}$ to $\bm{p}^{*}$.
2.2 Online Dual Algorithms
We introduce two basic dual-based algorithms for the OLP problem. We calculate and update the dual prices $\{\bm{p}_{t}\}_{t=1}^{T}$ and compute decisions $\{x_{t}\}_{t=1}^{T}$ based on rule (6).
2.2.1 LP-based Algorithm
In this method, we calculate the dual price by re-solving the online LP problem at each time step. The model is dynamically updated to incorporate real-time capacity constraints derived from the Action-history-dependent learning (AHDL) algorithm [16]. Define $d_{it}=\frac{b_{it}}{T-t}$ to be the average remaining resource capacity for resource $i$ at time $t$. The resulting optimization problem can be viewed as a $t$-sample approximation to the stochastic program specified by $\bm{d}_{t}=(d_{it},...,d_{mt})^{\top}$ as
$\min_{\bm{p}}$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ f_{\bm{d}_{t}}(\bm{p})\coloneqq\bm{d}_{t}^{\top}\bm{p}+\mathbb{E}[(r-\bm{a^{\top}p})^{+}]$ (7)
subject to $\leavevmode\nobreak\ \quad\bm{p}\geq\bm{0}.$
We denote the optimal solution to the stochastic program (7) as $\bm{p}_{t}^{*}$. Thus, instead of focusing on a static problem with fixed initial $\bm{d}$, we formulate a dynamic stochastic program that evolves according to the changing constraints. We outline this method in Algorithm 1.
Algorithm 1 LP-based Online AHDL algorithm
for $t$ = $1$ to $T$  do
       Observe $(r_{t},\bm{a}_{t})$ and make decision $x_{t}$ based on rule (6) if constraints are not violated. Update constraint for $i=1,...,m$:       remaining resource constraint $b_{it}=b_{i,t-1}-a_{it}x_{t}$       average resource capacity $d_{it}=\frac{b_{it}}{T-t}$ Solve the updated dual problem and obtain dual price $\bm{p}_{t+1}$:
$\bm{p}_{t+1}=\operatornamewithlimits{arg\,min}_{\bm{p}\geq\bm{0}}\ \bm{d}_{t}^{\top}\bm{p}+\frac{1}{t}\sum_{j=1}^{t}(r_{j}-\bm{a}_{j}^{\top}\bm{p})^{+}$
end for
This method incorporates the past decisions into the optimization of $\bm{p}_{t}$. If resources were over-utilized in earlier periods, the remaining supply decreases, prompting the algorithm to raise the dual price and become more selective with future orders. Conversely, if ample resources remain, the future dual price will be lowered, allowing more consumer requests to be accepted. This adaptive mechanism accounts for past actions by adjusting the available resource capacity.
2.2.2 Subgradient-based Algorithm
In this method, we calculate the dual price by iteratively moving in the direction of the steepest cost gradient. We maintain a static resource constraint, as the historical information is integrated through the gradient updates. Due to its lower computational overhead, the subgradient-based is well-suited for high-dimensional and real-time applications. Algorithm 2 summarizes this method, which is also known as the “first-order” method in [15].
Algorithm 2 First-Order Online algorithm
for $t$ = $1$ to $T$  do
       Observe $(r_{t},\bm{a}_{t})$ and make decision $x_{t}$ based on rule (6). Update learning rate $\alpha_{t}=\frac{1}{t+1}$. Compute subgradient and obtain dual price $\bm{p}_{t+1}$:
$\bm{p}_{t+1}=\operatornamewithlimits{arg\,min}_{\bm{p}\geq\bm{0}}\ (\bm{d}-\bm{a}_{t}x_{t})^{\top}\bm{p}+\frac{1}{2\alpha_{t}}\lVert\bm{p}-\bm{p}_{t}\rVert^{2}$
end for
We can derive a closed-form solution for the updated dual price from the above rule as
$\bm{p}_{t+1}$ $=(\bm{p}_{t}-\alpha_{t}(\bm{d}-\bm{a}_{t}x_{t}))\vee\bm{0}$
where the subgradient term evaluated at $\bm{p}_{t}$ is calculated as

### v3

# Wait-Less Offline Tuning and Re-solving for Online Decision Making

@@ Abstract
Online linear programming (OLP) has found broad applications in revenue management and resource allocation. State-of-the-art OLP algorithms achieve low regret by repeatedly solving linear programming (LP) subproblems that incorporate updated resource information. However, LP-based methods are computationally expensive and often inefficient for large-scale applications. By contrast, recent first-order OLP algorithms are more computationally efficient but typically suffer from weaker regret guarantees. To address these shortcomings, we propose a new algorithm that combines the strengths of LP-based and first-order OLP algorithms. Our algorithm re-solves the LP subproblems periodically at a predefined frequency $f$ and uses the latest dual prices to guide online decision-making. In parallel, a first-order method runs during each interval between LP re-solves and smooths resource consumption. Our algorithm achieves $\mathcal{O}(\log(T/f)+\sqrt{f})$ regret and delivers a “wait-less” online decision-making process that balances computational efficiency and regret guarantees. Extensive experiments demonstrate at least $10$-fold improvements in regret over first-order methods and $100$-fold improvements in runtime over LP-based methods.
Online Linear Programming, Sequential Decision-Making, Machine Learning
@@ 1 Introduction
Sequential decision-making has garnered significant attention for its utility in guiding optimal strategies in dynamic environments. The goal is to identify effective decisions and policies in environments where knowledge of the system continuously accumulates and evolves. Online Linear Programming (OLP) (Agrawal et al., 2014) offers a powerful framework that encapsulates the core principles of sequential decision-making and has been extensively applied to different domains, including resource allocation (Balseiro et al., 2022a), online advertising (Mehta et al., 2007), and inventory management (Talluri et al., 2004).
@@
We study an OLP problem where customers arrive sequentially, each requesting a combination of resources and offering a bidding price. The objective is to determine which resource requests to fulfill to maximize revenue while respecting resource constraints. The challenge is that decisions must be made immediately and irrevocably, relying solely on historical data without knowledge of future arrivals. The goal is to minimize regret with respect to the optimal hindsight linear programming (LP) solution.
@@
To guide real-time decision-making, state-of-the-art OLP algorithms estimate optimal dual prices and use them to make decisions. These algorithms fall into two main categories: LP-based and first-order methods. Specifically, LP-based methods update dual prices by repeatedly solving linear programs at each time step with all available information so far. However, the substantial computational demands limit their application in time-sensitive settings. First-order methods offer quick, incremental updates to dual prices using gradient information, but generally fall short of achieving the strong regret guarantee of LP-based methods. These trade-offs motivate an open question at the intersection of online learning and decision-making:
Can we simultaneously achieve
low regret and computational efficiency?
Our contributions.
We answer this question in the affirmative. We summarize our contributions as follows:
@@
  • Parallel Multi-Stage Framework: We separate online learning and decision-making into distinct processes that interact via a feedback loop. We re-solve the LP subproblems periodically at a fixed frequency and feed the updated dual price to guide decision-making until the next time that the LP is re-solved. To further enhance efficiency, we apply a first-order method during the initial and final stages, restarting with the most recent learning outcomes. By integrating the LP-based and first-order techniques, our approach leverages their respective strengths and achieves a balance between regret performance and computational costs.
@@
  • Regret Analysis: We unify the analysis of LP-based and first-order methods by deriving a new performance metric to account for their inter-dependency. Our analysis yields a “spectrum theorem” that bounds regret for any feasible choice of the re-solving frequency:
@@
    If we re-solve the LPs every $f$ time steps within horizon $T$ and apply a first-order method in the initial and final $f$ steps, we achieve a worst-case regret of $\mathcal{O}\big(\log(T/f)+\sqrt{f}\big).$
    In particular, $f=1$ yields a pure LP-based method with $\mathcal{O}(\log T)$ regret, while $f=T$ reduces to a pure first-order method with $\mathcal{O}(\sqrt{T})$ regret. By choosing an “optimal” $f$, one can achieve the best possible regret based on available computational resources and enable a “wait-less” decision-making system across all time steps.
  • Experiments: Through experiments across diverse distributions, we demonstrate that our algorithm achieves over a $10$-fold improvement in regret compared to first-order methods and over a $100$-fold improvement in runtime with comparable regret to LP-based methods.
Key Challenges.
OLP methods face a fundamental challenge in balancing efficient real-time decision-making and accurate dual-price learning. Specifically, LP-based methods ensure high-quality decisions with $\mathcal{O}(\log T)$ regret, but their computational costs increase quadratically with problem size. A natural alternative—batching customers and solving the LP once every $f$ arrivals—reduces this cost, but leaves customers waiting until the batch concludes. In contrast, first-order methods update dual prices via gradient information, allowing for faster computation. However, these methods require small step sizes to maintain decision quality, which slows their adaptation to new data. Even using different step sizes for learning and decision-making only achieves $\mathcal{O}(T^{1/3})$ regret when the distribution of customers has continuous support. Our parallel framework addresses these issues: an LP-based method periodically refines dual prices, while a first-order method immediately processes arriving customers using the most recent dual updates, thereby eliminating delays.
A second challenge emerges when analyzing the regret of this hybrid approach. LP-based methods rely on a stopping-time analysis (halting when resources are depleted), whereas first-order methods track constraint violations as part of the regret. Since these two formulations are fundamentally different, existing regret bounds cannot be naively combined. We overcome this challenge by introducing a unified performance metric that decomposes regret into three components: dual convergence, constraint violation, and leftover resources. This approach yields the first integrated analysis of LP-based and first-order OLP, culminating in our $\mathcal{O}(\log(T/f)+\sqrt{f})$ regret bound.
@@ 1.1 Related Literature
Online resource allocation and OLP problems have been widely studied under two predominant models: the stochastic input (Goel & Mehta, 2008; Devanur et al., 2019) and stochastic permutation models (Agrawal et al., 2014; Gupta & Molinaro, 2016). We study the former, where each customer’s resource request and bidding price are drawn i.i.d. from an unknown distribution. We use expected regret and constraint violation as the performance metric. We summarize some recent works in Table˜1.
LP-based Methods.
These methods solve the OLP dual repeatedly to update dual prices and make decisions. Early approaches enforced a fixed average resource constraint (Agrawal et al., 2014), whereas recent work dynamically tracks remaining resources and enables LP-based methods to achieve $\mathcal{O}(\log T)$ regret under continuous support and non-degeneracy assumptions (Li & Ye, 2022). Variants include multi-secretary problems (Bray, 2022), regularized resource constraints (Ma et al., 2024), and finite-support distributions yielding constant regret (Chen et al., 2024).
First-order Methods.
These methods generate decisions using gradient updates without solving LPs, enabling efficient computation. They achieve $\mathcal{O}(\sqrt{T})$ regret with mirror descent (Li et al., 2020; Balseiro et al., 2022a) and $\mathcal{O}(T^{3/8})$ under finite support and non-degeneracy assumptions (Sun et al., 2020). Variants include proximal updates (Gao et al., 2023), momentum-based mirror descent (Balseiro et al., 2022b), resource adjustments (Ma et al., 2024), and restart strategies yielding $\mathcal{O}(T^{1/3})$ regret (Gao et al., 2024).
Delay in Decision-making.
Table 1: Performances of Dual Algorithms in Recent OLP Literature
@@
(Li & Ye, 2022) Bounded, continuous support, non-degeneracy LP-based $\mathcal{O}(\log T\log\log T)$ Delay
(Bray, 2022) Bounded, continuous support, non-degeneracy LP-based $\mathcal{O}(\log T)$ Delay
(Chen et al., 2024) Bounded, finite support, non-degeneracy LP-based $\mathcal{O}(1)$ Delay
(Li et al., 2024) Bounded, finite support LP-based $\mathcal{O}(1)$ Delay
(Xu et al., 2024) Bounded, continuous support, non-degeneracy LP-based $\mathcal{O}(\log(T/f))$ Delay
@@
(Li et al., 2020) Bounded First-order $\mathcal{O}(\sqrt{T})$ No Delay
(Balseiro et al., 2022a) Bounded First-order $\mathcal{O}(\sqrt{T})$ No Delay
(Gao et al., 2023) Bounded First-order $\mathcal{O}(\sqrt{T})$ No Delay
(Sun et al., 2020) Bounded, finite support, non-degeneracy First-order $\mathcal{O}(T^{3/8})$ No Delay
(Gao et al., 2024) Bounded, continuous support, non-degeneracy First-order $\mathcal{O}(T^{1/3})$ No Delay
Delays arise from the time-consuming process of solving the large-scale, up-to-date LP subproblems for each customer. Golrezaei & Yao (2021) study a mix of impatient and partially patient customers, while Xie et al. (2023) show that batching requests can reduce regret. Concurrent work (Xu et al., 2024) reduces delay by solving LPs in batches but assumes a lower bound on resource requests and still requires waiting in initial and final batches. To achieve delay-free decisions, we shift the re-solving process offline and only fine-tune the solution online. Compared with previous works, our framework imposes standard assumptions, achieves $\mathcal{O}(\log(T/f)+\sqrt{f})$ regret, and ensures “wait-less” decision-making throughout the entire horizon.
Paper organization.
The rest of the paper is organized as follows. Section˜2 introduces the problem formulation and assumptions. Section 3 proposes our algorithms and main theoretical guarantee: a $\mathcal{O}(\log(T/f)+\sqrt{f})$ regret bound. Section 4 presents experiments to validate our theory.
@@ 2 Problem Setup
We use $\|\cdot\|$ to denote the Euclidean norm and $\langle\cdot\rangle$ to denote Euclidean inner product. Bold letters $\bm{A}$ and $\bm{a}$ denote matrices and vectors, respectively. Subscript $(\cdot)_{it}$ denotes the index for resource type $i$ at time $t$. The notation $(\cdot)^{+}=\max\{\cdot,0\}$ denotes the element-wise positive part function, and $\mathbb{I}(\cdot)$ denotes the 0-1 indicator function.
2.1 OLP Formulation
We study an online resource allocation problem over the time horizon $T$ under a stochastic input model. Initially, we have an inventory vector $\bm{b}\in\mathbb{R}^{m}$, representing $m$ resource types. At each time step $t$, a customer arrives with a request sampled i.i.d. as $(r_{t},\bm{a}_{t})\sim\mathcal{P}$, where $\bm{r}=(r_{1},\ldots,r_{T})^{\top}\in\mathbb{R}^{T}$ is the offered payment (bid), $\bm{A}=(\bm{a}_{1},\ldots,\bm{a}_{T})\in\mathbb{R}^{m\times T}$ is the matrix of customers’ resource demands, and $\mathcal{P}$ is a fixed, unknown distribution. We must decide whether to accept or reject each request, represented by the decision variables $\bm{x}=(x_{1},\ldots,x_{T})\in\{0,1\}^{T}$. The goal is to maximize the cumulative reward. This problem can be formulated as the following OLP, referred to as the primal linear program (PLP):
$\max_{\bm{0}\leq\bm{x}\leq\bm{1}}\langle\bm{r},\bm{x}\rangle\quad\text{s.t.}\quad\bm{Ax}\leq\bm{b}.$ (PLP)
@@ 2.1 OLP Formulation
$\min_{(\bm{p},\bm{y})\geq\bm{0}}\langle\bm{b},\bm{p}\rangle+\langle\bm{1},\bm{y}\rangle\quad\text{s.t.}\quad\bm{A}^{\top}\bm{p}+\bm{y}\geq\bm{r}$ (DLP)
where $\bm{p}$ is the vector of dual prices. Let $\bm{d}=\bm{b}/T\in\mathbb{R}^{m}$ denote the initial average resource. As demonstrated by (Li et al., 2020), (DLP) can be written as
$\min_{\bm{p}\geq\bm{0}}f_{T}(\bm{p})\coloneqq\tfrac{1}{T}\textstyle\sum_{t=1}^{T}\left[\langle\bm{d},\bm{p}\rangle+(r_{t}-\langle\bm{a}_{t},\bm{p}\rangle)^{+}\right]$ (1)
This formulation can be written as a $T$-sample approximation of the following stochastic program:
$\min_{\bm{p}\geq\bm{0}}f(\bm{p})\coloneqq\mathbb{E}[f_{T}(\bm{p})]=\bm{d}^{\top}\bm{p}+\mathbb{E}[(r-\bm{a^{\top}p})^{+}]$ (2)
where the expectation is taken with respect to $(r,\bm{a})\sim\mathcal{P}$.
We define optimal solutions to the $T$-sample approximation problem (1) and stochastic program (2) respectively as
$\bm{p}_{T}^{*}=\arg\min_{\bm{p}\geq\bm{0}}\penalty 10000\ f_{T}(\bm{p})\quad\text{and}\quad\bm{p}^{*}=\arg\min_{\bm{p}\geq\bm{0}}\penalty 10000\ f(\bm{p}).$
The decision variable $x_{t}^{*}$ can be established from the complementary slackness condition as
$x_{t}^{*}=\left\{\begin{array}[]{ll}\ 0,&r_{t}<\bm{a}_{t}^{\top}\bm{p}_{T}^{*},\vskip 2.0pt\\ \ 1,&r_{t}>\bm{a}_{t}^{\top}\bm{p}_{T}^{*}\end{array}\right$. (5)
@@
This connection between primal and dual solutions inspires OLP algorithms that make decisions based on dual prices:
$x_{t}=\mathbb{I}(r_{t}\geq\bm{a}_{t}^{\top}\bm{p}_{t}).$ (6)
2.2 Algorithms for OLP
We summarize two main dual-based OLP algorithms.
LP-based Method.
The LP-based method (Li & Ye, 2022) calculates dual prices by re-solving the online LP problem at each time step. Specifically, define $d_{it}=b_{it}/(T-t)$ as the average remaining resource for type $i$ at time $t$. The resulting optimization problem can be viewed as a $t$-sample approximation to the stochastic program specified by $\bm{d}_{t}=(d_{it},...,d_{mt})^{\top}$ as
$\min_{\bm{p}\geq\bm{0}}$ $\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ f_{\bm{d}_{t}}(\bm{p})\coloneqq\bm{d}_{t}^{\top}\bm{p}+\mathbb{E}[(r-\bm{a^{\top}p})^{+}]$ (7)
By updating $\bm{d}_{t}$, this method incorporates past decisions when computing $\bm{p}_{t}$. If resources are over-utilized in earlier periods, the supply decreases, prompting us to raise the dual price and be more selective with future orders. We outline this method in Algorithm˜3 of Section˜A.2.
First-order Method.
The first-order method (Li et al., 2020) calculates dual prices via online subgradient updates. We maintain a static average resource $\bm{d}=\bm{b}/T$, compute $x_{t}$ as per (6), and update the dual price as
$\bm{p}_{t+1}$ $=(\bm{p}_{t}-\alpha_{t}(\bm{d}-\bm{a}_{t}x_{t}))^{+}$
where the subgradient term evaluated at $\bm{p}_{t}$ is